# Concurrence fill and mode distribution of entanglement in neutrino oscillation

### v1

@@ Abstract
The flavor oscillations in neutrino system are known to be related to multi-mode entanglement of single particle state. The neutrino oscillations are shown to encompass not only quantum superposition among different mass eigenstates but also constitute a single particle multi-mode entanglement in the flavor basis as it propagates. In the framework of three flavor neutrino oscillation, we demonstrate that the measures of entanglement can be expressed in terms of experimentally accessible appearance and disappearance probabilities. We explicitly show here that the genuine tripartite entanglement measure, i.e., the tangle vanishes identically for all flavors signifying that three flavor neutrino system form a W-type entangled state. Further, we investigate alternative measures of tripartite entanglement like the partial tangle and the concurrence fill which capture the total sharing of entanglement beyond pairwise correlations. In terms of bipartite and bi-partitioned entanglement measures, we derive the symmetric invariant and the concurrence fill, which quantify the distributed entanglement completely expressible in terms of flavor transition probabilities. These entanglement measures display distinct energy dependent patterns across the oscillation window which can be experimentally accessible in the long baseline experiments like DUNE providing an alternative quantum information perspective on flavor evolution. We use GLobal Long Baseline Experiment Simulator (GLoBES) simulations within the DUNE set up to investigate these tripartite entanglement measures in terms of neutrino energy and the length of the baseline. It is observed that, at the point of maximal mixing, these measures show near maximal entanglement between the muon and the tau flavor modes establishing entanglement monogamy. Within the DUNE set up, the wide band of energy and expected higher sensitivity to CP-violation at second oscillation maximum provide a unique advantage to explore the quantum correlation effects across a broader energy window. Thus, the flavor coherence and CP-phase dependent interference can influence the quantum entanglement measures among neutrino flavor modes and thereby, such potential new physics effects can be probed in current or future planned neutrino experiments.
@@ 1 Introduction
Neutrino oscillations provide a quantum mechanical description of flavor transitions among the three neutrino flavors $(\nu_{e},\nu_{\mu},\nu_{\tau})$. In the standard three-flavor picture, a neutrino flavor state $\nu_{\alpha}$ $(\alpha=e,\mu,\tau)$ is expressed as a quantum superposition of the three mass eigenstates $\nu_{i}$ $(i=1,2,3)$. A neutrino is created in a definite flavor state. As the neutrino propagets, each mass eigenstate accumulates a different phase depending on the baseline and energy, and the interference among these phases gives rise to the oscillation among differnt flavors over distance. This interference of mass components modulates the flavor content and generates the oscillation phenomena observed in accelerator, atmospheric, solar, and reactor experiments Bilenky:2004xm ; Kayser:2012ce ; SNO:2002tuh ; T2K:2011ypd ; MINOS:2011amj ; K2K:2006yov ; DayaBay:2012fng ; Super-Kamiokande:2002ujc ; NOvA:2016vij . The oscillation pattern is governed by the mass-squared differences $(\Delta m^{2}_{21},|\Delta m^{2}_{31}|)$, three mixing angles $(\theta_{12},\theta_{13},\theta_{23})$, the leptonic CP violating phase $\delta$, and the dynamical quantities $L$, $E$, and the matter potential $A=2\sqrt{2}\,G_{F}N_{e}E$.
@@
Quantum entanglement, on the other hand, is a fundamental nonclassical feature that arises directly from the superposition principle and plays a central role in quantum information science Nielsen_Chuang_2010 ; PhysRevA.98.052351 . In controlled quantum systems, entanglement serves as a valuable resource that makes possible a variety of remarkable protocols—including quantum teleportation and quantum key distribution PhysRevLett.70.1895 ; PhysRevLett.67.661 . Neutrino oscillations and quantum entanglement both arise from the principle of superposition, but they play fundamentally different roles. In oscillations, superposition governs how a single neutrino evolves among flavor states, while in entanglement it describes inseparable correlations shared between different subsystems. Superposition concerns coherent combinations within a single subsystem, while entanglement refers to non factorizable correlations between distinct subsystems.
@@
Nevertheless, the three-mode structure of neutrino oscillations naturally raises the question of whether the evolving flavor state should be interpreted purely as a coherent superposition or as a form of single-particle multimode entanglement. This possibility has been explored in a number of works within the occupation-number representation, where the three flavor (or mass) modes define an effective tripartite system Blasone:2007vw ; Blasone:2010ta ; Blasone:2013zaa ; Blasone:2014jea ; Blasone:2015lya ; Banerjee:2015mha ; Ettefaghi:2023zsh ; Blasone:2022joq ; Li:2022mus ; Patwardhan:2021rej ; Ettefaghi:2020otb . However, a complete and experimentally relevant characterization of this single-particle multipartite entanglement, particularly in the context of realistic long baseline experiments, remains largely unexplored. This challenge is compounded by the fact that a neutrino continuously oscillates among separable and maximally entangled configurations as it evolves Banerjee:2024lih .
@@
In the context of three flavor neutrino system early theoretical studies demonstrated that flavor states, when formulated in the occupation number (mode) representation, intrinsically possess entanglement because the flavor content at any time is a coherent admixture of mass components Blasone:2007vw ; Blasone:2010ta ; Blasone:2013zaa ; Blasone:2014jea ; Blasone:2015lya ; Banerjee:2015mha ; Jha:2020dav ; Jha:2022yik ; Cirigliano:2024pnm . Follow up studies have shown that, once the three flavor states are treated within a common Hilbert space, both bipartite and tripartite forms of entanglement can arise during neutrino propagation. These analyses demonstrated that flavor oscillations continuously reorganize the underlying quantum correlations among the three modes, mirroring the correlation dynamics known from multimode quantum-information systems Ettefaghi:2023zsh ; Blasone:2022joq ; Li:2022mus ; Patwardhan:2021rej ; Ettefaghi:2020otb . More recent investigations have extended this perspective to phenomenologically relevant settings, using entanglement based diagnostics to probe oscillation parameters, matter effects, and possible nonstandard interactions (NSI) Banerjee:2024lih ; Konwar:2024nrd ; Konwar:2025ipv . Complementary work on the quantum speed limit has further clarified the dynamical constraints governing flavor evolution Bouri:2024kcl ; Jha:2025ekn . Together, these developments highlight that neutrino oscillations can be understood not only as flavor transitions but also as the evolution of a single-particle, multimode entangled state shaped by mixing dynamics and propagation effects.
However entanglement quantification for multipartite systems is considerably more intricate than for bipartite ones Mukherjee:2025edo . Well established bipartite measures such as concurrence Wootters:1997id ; Hill:1997pfa , negativity Vidal:2002zz , and entanglement of formation Wootters:1997id , do not trivially generalize to three mode systems. For tripartite qubit systems, the two inequivalent classes of genuine multipartite entanglement, GHZ and W states, are distinguished by the tangle (or three tangle) Coffman:1999jd , which detects GHZ-type correlations but vanishes for W states. Additional tools such as partial tangles Lee:2005qzt help diagnose biseparability. More recently, the concurrence fill $C_{F}$ Li:2021epj has been proposed as a geometric measure of genuine multipartite entanglement, based on the area enclosed by the squared concurrences Sarkar:2024myo . This measure satisfies the essential axioms of non negativity, monotonicity, discriminance, convexity, and smoothness, and is particularly sensitive to distributed W class of entanglement PhysRevLett.127.040403 ; Aggarwal:2025fab .
@@
Motivated by this growing theoretical foundation, in this work we provide a comprehensive characterization of bipartite and genuine tripartite entanglement in the three-flavor neutrino system, expressing all measures directly in terms of experimentally measurable oscillation probabilities. A core objective of this work is to determine and to characterize in an experimentally meaningful way the specific class of quantum entanglement that emerges in the evolution of three flavor neutrino oscillations. We investigate the multipartite entanglement structure of the three flavor neutrino system using experimentally accessible quantities, i.e., appearance and disappearance probabilities in Deep Underground Neutrino Experiment (DUNE) thereby establishing a direct bridge between quantum information theory and long baseline neutrino phenomenology. We first construct a single particle, multimode description of an oscillating neutrino and show how the standard oscillation formalism emerges naturally from the entangled flavor state. We then introduce three complementary entanglement diagnostics tailored to the neutrino system: (i) tangle, (ii) partial tangles, and (iii) the concurrence fill. A key aspect of this work is that we provide explicit expressions for these measures directly in terms of neutrino oscillation amplitudes and measurable probabilities. We further perform detailed DUNE-based simulations to map these quantifiers as functions of energy and baseline, thereby identifying the dynamically accessible entanglement regimes.
@@
Our results show that the tangle vanishes across the entire energy range, ruling out GHZ-type entanglement and suggesting separable, biseparable, or W-type structure. Computation of the pairwise concurrences reveals that entanglement is dominantly shared within the $\nu_{\mu}-\nu_{\tau}$ and $\nu_{e}-\nu_{\tau}$ pairs, with negligible correlations in the $\nu_{\mu}-\nu_{e}$ sector. Together with the partial-tangle analysis, this pattern provides clear evidence of biseparability in the tripartite neutrino state, both in matter and in vacuum. To further assess genuine tripartite structure, we evaluate the concurrence fill and find that the area of the concurrence triangle varies with energy. A nonzero area indicates genuine three-mode entanglement, consistent with a W-type configuration, while the area vanishes at specific energies where tripartite correlations disappear. Analysis of the W-class inequality confirms that the neutrino system indeed occupies regions of W-type entanglement over significant portions of its evolution.
@@
The structure of the paper is as follows. Section 2 establishes the theoretical formalism and the mode-distribution picture of entanglement in neutrino oscillations. In Section 3, we derive the multipartite entanglement measures—tangle, partial tangle, and concurrence fill—expressed directly in terms of reduced density matrices and oscillation probabilities. Section 4 details the DUNE simulation framework and presents numerical results, quantifying W-type entanglement as a function of energy and baseline under various mixing scenarios. Finally, Section 5 summarizes our findings and discusses experimental prospects and future directions.
@@ 2 Theoretical Framework
The effective Hamiltonian in the flavor basis, for three flavor neutrino oscillation in presence of matter, is expressed as Pontecorvo:1957qd ; Maki:1962mu ; Pontecorvo:1967fh ; Patra:2023ltl ,
@@
where $\alpha=\Delta m_{21}^{2}/\Delta m_{31}^{2}$ characterizes the ratio between solar and atmospheric mass square differences. The matter effect are encoded in $\hat{A}=A/\Delta m^{2}_{31}$ with $A=2V_{CC}$. The unitray mixing matrix diagonalizing the light neutrino masses turns out to be the same as well known Pontecorvo, Maki, Nakagawa, Sakata (PMNS) mixing matrix, i.e, $U_{\rm PMNS}\equiv U$ in the basis where the charged leptons are already diagonal. Using the standard parametrization in terms of three neutrino mixing angles $\theta_{ij}$ and a Dirac CP phase $\delta_{\rm CP}$, the form of neutrino mixing matrix $U$ is given by
@@
For the case of three flavors in vaccum, the neutrino flavor states are defined as
@@
Using Eq.(2), the time evolution relation for the mass eigen states and reverting back to the flavor basis, the evolution operator for the flavor basis $\widetilde{U}\big(t,\theta,\delta_{\rm CP}\big)$ is given by
@@
In contrast to conventional two-particle entanglement, where correlations arise between physically distinct subsystems, single-particle entanglement (SPE)  [^1] The single particle multi-mode entanglement has already been investigated in Ref. PhysRevA.72.064306 . The experimental verification along with the non-local correlation of single particle entanglement has also been established by Pasini et. al. PhysRevA.102.063708 in photon system. refers to the distribution of quantum superposition across different modes of a single quantum system. In the neutrino context, the three flavor modes $(\nu_{e},\,\nu_{\mu},\,\nu_{\tau})$ naturally define three orthogonal Hilbert-space sectors, each of which can be treated as a three level system in the occupation number basis,
@@
Tracing out mode subspaces reduces the full $8\times 8$ density matrix to an effective $3\times 3$ object. This reduced description aligns naturally with the standard three-flavor oscillation formalism and provides a clean algebraic framework for analyzing coherence and entanglement distribution among $\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$. Therefore within this single excitation subspace, the time evolved flavor state takes the form

### v2

@@ Abstract
The flavor oscillations in the neutrino system are known to be related to the multi-mode entanglement of a single particle state. Neutrino oscillations are shown to encompass not only quantum superposition among different mass eigenstates but also a single-particle multi-mode entanglement in the flavor basis as they propagate. In the framework of three flavor neutrino oscillations, we demonstrate that the measures of entanglement can be expressed in terms of experimentally accessible appearance and disappearance probabilities. We explicitly show here that the genuine tripartite entanglement measure, i.e., the tangle, vanishes identically for all flavors, signifying that a three flavor neutrino system forms a W-type entangled state. Further, we investigate alternative measures of tripartite entanglement like the partial tangle and the concurrence fill, which capture the total sharing of entanglement beyond pairwise correlations. In terms of bipartite and bipartitioned entanglement measures, we derive the symmetric invariant and the concurrence fill, which quantify the distributed entanglement and are completely expressible in terms of flavor transition probabilities. These entanglement measures display distinct energy dependent patterns across the oscillation window, which can be experimentally accessible in the long baseline experiments like DUNE, providing an alternative quantum information perspective on flavor evolution. We use General Long Baseline Experiment Simulator (GLoBES) simulations within the DUNE setup to investigate these tripartite entanglement measures in terms of neutrino energy and the length of the baseline. It is observed that, at the point of maximal mixing, these measures show near maximal entanglement between the muon and the tau flavor modes, establishing entanglement monogamy. Within the DUNE setup, the wide band of energy and expected higher sensitivity to CP violation at the second oscillation maximum provide a unique advantage to explore the quantum correlation effects across a broader energy window. Thus, the flavor coherence and CP-phase dependent interference can influence the quantum entanglement measures among neutrino flavor modes, and thereby, such potential new physics effects can be probed in current or future planned neutrino experiments.
@@ 1 Introduction
Neutrino oscillations provide a quantum mechanical description of flavor transitions among the three neutrino flavors $(\nu_{e},\nu_{\mu},\nu_{\tau})$. In the standard three-flavor picture, a neutrino flavor state $\nu_{\alpha}$ $(\alpha=e,\mu,\tau)$ is expressed as a quantum superposition of the three mass eigenstates $\nu_{i}$ $(i=1,2,3)$. A neutrino is created in a definite flavor state. As the neutrino propagates, each mass eigenstate accumulates a different phase depending on the baseline and energy, and the interference among these phases gives rise to oscillations among different flavors over distance. This interference of mass components modulates the flavor content and generates the oscillation phenomena observed in accelerator, atmospheric, solar, and reactor experiments Bilenky:2004xm ; Kayser:2012ce ; SNO:2002tuh ; T2K:2011ypd ; MINOS:2011amj ; K2K:2006yov ; DayaBay:2012fng ; Super-Kamiokande:2002ujc ; NOvA:2016vij . The oscillation pattern is governed by the mass-squared differences $(\Delta m^{2}_{21},|\Delta m^{2}_{31}|)$, three mixing angles $(\theta_{12},\theta_{13},\theta_{23})$, the leptonic CP violating phase $\delta$, and the dynamical quantities $L$, $E$, and the matter potential $A=2\sqrt{2}\,G_{F}N_{e}E$.
@@
Quantum entanglement, on the other hand, is a fundamental nonclassical feature that arises directly from the superposition principle and plays a central role in quantum information science Nielsen_Chuang_2010 ; PhysRevA.98.052351 . In controlled quantum systems, entanglement serves as a valuable resource that makes possible a variety of remarkable protocols—including quantum teleportation and quantum key distribution PhysRevLett.70.1895 ; PhysRevLett.67.661 . Neutrino oscillations and quantum entanglement both arise from the principle of superposition, but they play fundamentally different roles. In oscillations, superposition governs how a single neutrino evolves among flavor states, while in entanglement, it describes inseparable correlations shared between different subsystems. Superposition concerns coherent combinations within a single subsystem, while entanglement refers to non factorizable correlations between distinct subsystems.
@@
Nevertheless, the three-mode structure of neutrino oscillations naturally raises the question of whether the evolving flavor state should be interpreted purely as a coherent superposition or as a form of single-particle multimode entanglement. This possibility has been explored in several works within the occupation-number representation, where the three flavor (or mass) modes define an effective tripartite system Blasone:2007vw ; Blasone:2010ta ; Blasone:2013zaa ; Blasone:2014jea ; Blasone:2015lya ; Banerjee:2015mha ; Ettefaghi:2023zsh ; Blasone:2022joq ; Li:2022mus ; Patwardhan:2021rej ; Ettefaghi:2020otb . However, a complete and experimentally relevant characterization of this single-particle multipartite entanglement, particularly in the context of realistic long baseline experiments, remains largely unexplored. This challenge is compounded by the fact that a neutrino continuously oscillates among separable and maximally entangled configurations as it evolves Banerjee:2024lih .
@@
In the context of three flavor neutrino system, early theoretical studies demonstrated that flavor states, when formulated in the occupation number (mode) representation, intrinsically possess entanglement because the flavor content at any time is a coherent admixture of mass components Blasone:2007vw ; Blasone:2010ta ; Blasone:2013zaa ; Blasone:2014jea ; Blasone:2015lya ; Banerjee:2015mha ; Jha:2020dav ; Jha:2022yik ; Cirigliano:2024pnm . Follow up studies have shown that, once the three flavor states are treated within a common Hilbert space, both bipartite and tripartite forms of entanglement can arise during neutrino propagation. These analyses demonstrated that flavor oscillations continuously reorganize the underlying quantum correlations among the three modes, mirroring the correlation dynamics known from multimode quantum-information systems Ettefaghi:2023zsh ; Blasone:2022joq ; Li:2022mus ; Patwardhan:2021rej ; Ettefaghi:2020otb .
More recent investigations have extended this perspective to phenomenologically relevant settings, using entanglement based diagnostics to probe oscillation parameters, matter effects, and possible nonstandard interactions (NSI) Banerjee:2024lih ; Konwar:2024nrd ; Konwar:2025ipv . Complementary work on the quantum speed limit has further clarified the dynamical constraints governing flavor evolution Bouri:2024kcl ; Jha:2025ekn ; Jha:2026jef . In addition, the role of entanglement among supernova neutrinos and its potential impact on observable signals at the DUNE experiment have been investigated in Ref. Johny:2026obb . It may further be noted that quantum entanglement in the neutrino system has been studied from various complementary perspectives. In particular, in Ref. Bittencourt:2023asd , an analysis of tripartite entanglement in three flavor neutrino oscillation was carried out using complete complementary relations. It was explicitly demonstrated that, unlike pure bipartite neutrino states with vanishing local coherence, tripartite neutrino systems generally exhibit non-vanishing local coherences. Extending these ideas, subsequently, a triality relation for three-flavor neutrino oscillations was explored in terms of predictability, visibility, and I-concurrence Banerjee:2025vyh . In parallel, a systematic study of quantum coherence in both two- and three-flavor neutrino oscillations has been performed, examining its energy and baseline dependence within the standard oscillation framework. This analysis complements the present entanglement results by isolating the role of coherence in governing flavor interference and CP-phase effects Alok:2025qqr .
Together, these developments highlight that neutrino oscillations can be understood not only as flavor transitions but also as the evolution of a single-particle, multimode entangled state shaped by mixing dynamics and propagation effects.
[…]
@@
Motivated by this growing theoretical foundation, in this work, we provide a comprehensive characterization of bipartite and genuine tripartite entanglement in the three-flavor neutrino system, expressing all measures directly in terms of experimentally measurable oscillation probabilities. A core objective of this work is to determine and to characterize in an experimentally meaningful way the specific class of quantum entanglement that emerges in the evolution of three flavor neutrino oscillations. Therefore, to make the global entanglement structure transparent, we have considered the W and GHZ states, which differ clearly in their physical properties. We investigate the multipartite entanglement structure of the three flavor neutrino system using experimentally accessible quantities, i.e., appearance and disappearance probabilities in Deep Underground Neutrino Experiment (DUNE), thereby establishing a direct bridge between quantum information theory and long baseline neutrino phenomenology. We first construct a single particle, multimode description of an oscillating neutrino and show how the standard oscillation formalism emerges naturally from the entangled flavor state. We then introduce three complementary entanglement diagnostics tailored to the neutrino system: (i) tangle, (ii) partial tangles, and (iii) the concurrence fill. A key aspect of this work is that we provide explicit expressions for these measures directly in terms of neutrino oscillation amplitudes and measurable probabilities. We further perform detailed DUNE-based simulations to map these quantifiers as functions of energy and baseline, thereby identifying the dynamically accessible entanglement regimes.
@@
Our results show that the tangle vanishes across the entire energy range, ruling out GHZ-type entanglement and suggesting separable, biseparable, or W-type structure. Computation of the pairwise concurrences reveals that entanglement is dominantly shared within the $\nu_{\mu}-\nu_{\tau}$ and $\nu_{e}-\nu_{\tau}$ pairs, with negligible correlations in the $\nu_{\mu}-\nu_{e}$ sector. Together with the partial-tangle analysis, this pattern provides clear evidence of biseparability in the tripartite neutrino state, both in matter and in vacuum. To further assess the genuine tripartite structure, we evaluate the concurrence fill and find that the area of the concurrence triangle varies with energy. A nonzero area indicates genuine three-mode entanglement, consistent with a W-type configuration, while the area vanishes at specific energies where tripartite correlations disappear. Analysis of the W-class inequality confirms that the neutrino system indeed occupies regions of W-type entanglement over significant portions of its evolution.
In addition to characterizing the structure of flavor-mode entanglement, it is also important to examine whether such quantum correlations can provide practical insight into the fundamental parameters of the oscillation framework. In particular, the sensitivity to the Dirac CP phase $\delta_{CP}$ is one of the central goals of long-baseline experiments. Since the entanglement measures considered in this work are directly expressible in terms of oscillation probabilities, their variation with $\delta_{CP}$ carries potential information about CP violation. This motivates us to investigate not only the magnitude of entanglement but also its phase sensitivity in order to assess whether regions of enhanced entanglement response can serve as useful indicators for optimizing experimental measurements. Such a study of the correlation between $\delta_{\rm CP}$ and the entanglement measures is of particular relevance for the DUNE experiment, which covers a wide range of energies.
@@
The structure of the paper is as follows. Section 2 establishes the theoretical formalism and the mode-distribution picture of entanglement in neutrino oscillations. In Section 3, we derive the multipartite entanglement measures, tangle, partial tangle, and concurrence fill, expressed directly in terms of reduced density matrices and oscillation probabilities. Section 4 details the DUNE simulation framework and presents numerical results, quantifying W-type entanglement as a function of energy and baseline under various mixing. In this section, we further analyze the sensitivity to the Dirac CP phase, focusing on the phase dependence of the concurrence fill $C_{F}$ as a potential probe of CP violation. Finally, Section 5 summarizes our findings and discusses experimental prospects and future directions.
@@ 2 Theoretical Framework
The effective Hamiltonian in the flavor basis, for three flavor neutrino oscillation in the presence of matter, is expressed as Pontecorvo:1957qd ; Maki:1962mu ; Pontecorvo:1967fh ; Patra:2023ltl ,
@@
where $\alpha=\Delta m_{21}^{2}/\Delta m_{31}^{2}$ characterizes the ratio between solar and atmospheric mass square differences. The matter effect are encoded in $\hat{A}=A/\Delta m^{2}_{31}$ with $A=2V_{CC}$. The unitary mixing matrix diagonalizing the light neutrino masses turns out to be the same as well known Pontecorvo, Maki, Nakagawa, Sakata (PMNS) mixing matrix, i.e, $U_{\rm PMNS}\equiv U$ on the basis where the charged leptons are already diagonal. Using the standard parametrization in terms of three neutrino mixing angles $\theta_{ij}$ and a Dirac CP phase $\delta_{\rm CP}$, the form of the neutrino mixing matrix $U$ is given by
@@
For the case of three flavors in vacuum, the neutrino flavor states are defined as
@@
Using Eq.(2), the time evolution relation for the mass eigen states and reverting to the flavor basis, the evolution operator for the flavor basis $\widetilde{U}\big(t,\theta,\delta_{\rm CP}\big)$ is given by
@@
In contrast to conventional two-particle entanglement, where correlations arise between physically distinct subsystems, single-particle entanglement (SPE)  [^1] The single particle multi-mode entanglement has already been investigated in Ref. PhysRevA.72.064306 . The experimental verification along with the non-local correlation of single particle entanglement has also been established by Pasini et. al. PhysRevA.102.063708 in a photon system. refers to the distribution of quantum superposition across different modes of a single quantum system. In the neutrino context, the three flavor modes $(\nu_{e},\,\nu_{\mu},\,\nu_{\tau})$ naturally define three orthogonal Hilbert-space sectors, each of which can be treated as a three level system in the occupation number basis,
@@
Tracing out mode subspaces reduces the full $8\times 8$ density matrix to an effective $3\times 3$ object. This reduced description naturally aligns with the standard three-flavor oscillation formalism and provides a clean algebraic framework for analyzing the coherence and entanglement distribution among $\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$. Therefore, within this single excitation subspace, the time-evolved flavor state takes the form